Help me please

Substance

S° (J/mol×K)

Al (s)

28.33

Al2O3 (s)

50.92

Cl2 (g)

223.07

H2 (g)

130.68

N2 (g)

191.61

O2 (g)

205.14

Si (s)

18.83

SiCl4 (g)

330.73

SiHCl3 (g)

330.73

SiH2Cl2 (g)

285.71

SiH3Cl (g)

313.87

SiH4 (g)

204.62

Si3N4 (s)

101.3

SiO2 (s)

41.84

1. Predict the sign of \Delta S°rxn for the following processes. Then calculate \Delta S°rxn using the standard values just given.

sign of \Delta S°rxn

\DeltaS°rxn

a. SiO2 (s) \rightarrow Si (s) + O2 (g)

b. Al2O3 (s) \rightarrow 2 Al (s) + 3/2 O2 (g)

c. Si (s) + 2 Cl2 (g) \rightarrow SiCl4 (g)

d.3 Si (s) + 2 N2 (g) \rightarrow Si3N4 (s)

Thank you!
Answer #1

sign of \DeltaS°rxn

\DeltaS°rxn

a. SiO2 (s) \rightarrow Si (s) + O2 (g)

+ ve (positive)

b. Al2O3 (s) \rightarrow 2 Al (s) + 3/2 O2 (g)

+ ve (positive)

c. Si (s) + 2 Cl2 (g) \rightarrow SiCl4 (g)

- ve (negative)

d.3 Si (s) + 2 N2 (g) \rightarrow Si3N4 (s)

- ve (negative)

Explanation :

a) SiO2 (s) \rightarrow Si (s) + O2 (g)

\DeltaS°rxn = ( 18.83 + 205.14 ) - 41.84

\DeltaS°rxn   = 182.13 J/mol K

b) b. Al2O3 (s) \rightarrow 2 Al (s) + 3/2 O2 (g)

\DeltaS°rxn = (2 x 28.33 + 3/2 x 205.14) - 50.92

             = 313.45 J/mol . K

c)

\DeltaS°rxn = 330.73 - (18.83 + 2 x 223.07)

              = - 134.24 J/mol K

d)

\DeltaS° rxn = 101.3 - 18.83 +2 x 191.61

               = - 300.75 J/mol K